# On the Origin of Kinematic Structure in the Young Association Serpens OB2

### v1

@@ Abstract
The Serpens OB2 association ($\ell\sim 18\aas@@fstack{\circ}5$, $b\sim 1\aas@@fstack{\circ}9$, $d=1950\pm 30$ pc) is a large star-forming complex $\sim$65 pc above the Galactic midplane, with a clumpy, elongated structure extending $\sim$50 pc parallel to the plane. We analyse a sample of probable association members, including OB stars and low-to-intermediate-mass young stellar objects (YSOs) from the SPICY catalogue. While both populations are found throughout the association, the OB stars lie preferentially on the side nearest the Galactic plane, while the YSOs are generally younger and more strongly clustered around molecular-cloud clumps detected in 13CO MWISP data. Using Gaia DR3 proper motions to probe the association’s internal kinematics, we find aligned stellar velocities on length scales $\lesssim$2 pc, two-point velocity statistics that show increasing velocity differences and predominantly divergent motions at larger separations, and distinct velocities for star clusters within the association. Finally, the association exhibits gradual, but statistically significant global expansion perpendicular to the Galactic plane, with a spatial gradient of $0.10\pm 0.02$ km s-1 pc-1. The expansion of the H ii region Sh 2-54, powered by the association’s OB stars, may be accelerating the star-forming cloud away from the plane given the system’s geometry, plausibly inducing the vertical stellar velocity gradient. The clumpy stellar distribution, correlated velocities on small scales, and increasingly divergent motions on larger scales are consistent with an initial velocity field inherited from a turbulent molecular cloud modified by stellar feedback. Ser OB2 demonstrates that the multi-scale expansion of an OB association can begin even while star formation is still ongoing throughout the complex.
@@ 3.2 Member Properties
The $J$ versus JHJ-H colour–magnitude diagram is shown in Fig. 3. Using PARSEC stellar evolution models (Bressan et al., 2012) with an age of 4.4 Myr (Kharchenko et al., 2005), an extinction of $A_{V}\sim 4$ mag reproduces the OB locus but lies significantly bluer than most YSOs. For the 4.4 Myr isochrone to pass through the middle of the YSO distribution on the diagram, it would need to be reddened by $A_{V}\sim 10$ mag. Such an extinction is implausibly high given that many of these stars have Gaia astrometry. For example, a 2 M star with this age and extinction would have a Gaia magnitude of $G=20.6$ mag, which is too faint for reliable astrometry, yet dereddening on the near-infrared diagram suggests that some YSOs are in this mass range. A younger isochrone alleviates this tension. With a 1 Myr isochrone, an extinction of $A_{V}\sim 6$ mag is sufficient to explain the YSO colours, and a 2 M star would then have $G=18.4$ mag, within the range where Gaia provides astrometry. Owing to differential extinction in massive star-forming regions and degeneracies between effective temperature and extinction in spectral-energy-distribution fitting, it is difficult to obtain precise masses and ages for individual stars. Nevertheless, the constraints provided by the colour-magnitude diagram suggest that the YSOs are consistent with an age of $\sim$1 Myr and they are likely to be, on average younger and more extincted than the OB population.
@@ 3.3 Extrapolated Total Stellar Membership
Assuming our OB member list is complete and no evidence of past supernovae, we can extrapolate the total stellar population of the association by scaling from the initial mass function (IMF). The Maschberger (2013) IMF predicts that OB stars ($M>8$ $M_{\odot}$) comprise $\sim$0.7% of all newly formed stellar-mass objects ($M>0.08$ $M_{\odot}$). Given $\sim$80 probable OB stars in Ser OB2 after removing those with inconsistent parallaxes, the association is likely to have a total population of $\sim\!\!10^{4}$ stellar-mass members. For comparison, Luhman (2022) estimates 8300 stellar-mass members in Sco–Cen, meaning that these two associations have similar stellar population numbers.
@@ 6 Molecular Cloud
In the centre of the association, a channel, empty of molecular gas, extends vertically (in Galactic coordinates) through the cloud. The channel is also visible as a vertical gap in the nebulosity in the Spitzer/IRAC images (Fig. 1), with walls traced by polycyclic aromatic hydrocarbon (PAH) emission. Several OB stars lie in projection within this channel, suggesting they helped ionise the H ii region and clear a passage through the cloud. Because many of the OB stars (including those in NGC 6604) lie below the cloud in Galactic latitude while the thermal chimney extends upward, this channel may have served as the pathway where the H ii region broke through the cloud and inflated the chimney.
Figure 10: The MWISP 13CO J=1J=1–0 zeroth moment map between 25 and 35 km s-1. The greyscale image shows integrated emission in units of $\mathrm{K}\cdot\mathrm{km}$ s-1. OB stars are marked by blue crosses, Ser OB2 clusters (Table 2) are labelled, and the stellar surface density contours from Fig. 4 are drawn in yellow.
@@ 7 Discussion
We observe distinct velocity patterns on different spatial scales. On the smallest scales probed by our sample ($\lesssim 2$ pc), the $\theta_{ij}$ statistic indicates that nearby stars have preferentially aligned velocity directions. These scales encompass typical sizes of individual young clusters and subclusters (Lada & Lada, 2003; Kuhn et al., 2014, 2017), whose stars may have inherited similar initial velocities from forming in the same molecular clump. The VSAT statistics probe pairwise velocity differences over more than an order of magnitude in scale. For separations of $\sim$2–20 pc, both statistics show increasing velocity differences with separation, reminiscent of the Larson (1981) relations where cloud velocity dispersions increase with length scale. The VSAT $\Delta v_{ij,D}$ statistic has a positive median for separations >>2 pc, indicating net divergent motions, with divergence increasing with spatial separation. On intermediate scales, several clusters are moving outward, but others deviate from this trend, implying a more complex velocity field. On the largest scales, we detect gradual global expansion of the association along the axis perpendicular to the Galactic plane.
A possible scenario to explain the asymmetric global expansion is sequential star formation in a cloud accelerating away from the plane. Previous work has shown that the $\sim$4–5 Myr-old OB population in Ser OB2 is energetically capable of driving the large H ii region Sh 2-54 and the associated ‘Stockert chimney’ (Kundt & Mueller, 1987; Forbes, 2000). [^6] An expanding bubble model has been used to successfully describe another thermal chimney in the W4 region, obtaining an expansion age of 2.5 Myr (Basu et al., 1999). While both chimneys are of similar length, the Ser OB2 chimney is much narrower and its ionising source is less certain (Normandeau & Basu, 1999). Although the interface between the Sh 2-54 H ii region and the molecular cloud is likely to be complex (Section 6), the OB stars are preferentially nearer the Galactic plane, so the net force of the expanding H ii region is likely to be directed upward (i.e., in the $\eta$ direction). If the expansion of the Sh 2-54 H ii region has been pushing the cloud away from the midplane, stars forming in this accelerated gas would inherit upward velocities, producing a vertical velocity gradient.
@@ 8 Conclusion
Ser OB2 is a large OB association notable for its youth ($\lesssim 5$ Myr), its moderate height above the Galactic midplane (65 pc), and its role in launching a $\sim$200 pc thermal chimney. We have analysed a sample of Ser OB2 members assembled from OB stars in the ALS III catalogue and YSOs from the SPICY catalogue, with membership vetted using Gaia DR3 parallaxes. Owing to the inclusion of YSOs, this is the first study of the region to incorporate a substantial sample of low- and intermediate-mass members. The OB stars and YSOs have broadly similar spatial distributions and kinematics, but the OB stars lie closer to the Galactic midplane while the YSOs cluster more tightly around molecular-cloud clumps.
@@
  • While previous work indicates an age of 4–5 Myr for the OB population (e.g. Forbes, 2000), the same age would imply implausibly large extinctions for the low-mass YSOs. The YSOs are more consistent with an age of $\sim$1 Myr, implying an overall age spread of a few Myr in the association (Section 3.2).
  • The Ser OB2 members exhibit a clumpy spatial distribution elongated parallel to the Galactic midplane. We identified six discrete stellar clusters (NGC 6604, Gum 85, Dobashi 782, 815, 822, and G18.7+1.8) superimposed on a relatively smooth, centrally concentrated population (Section 4). Five of these clusters are associated with distinct molecular cloud clumps in 13CO J=1J=1–0 MWISP data cubes. Although the spatial distributions of OB stars and YSOs overlap, the OB stars are preferentially closer to the Galactic midplane.
@@
The disruption of clusters within Ser OB2 may contribute to the region’s overall expansion, although it remains difficult to assess from our data whether the clusters detected in this region are gravitationally bound given sparse sampling and Gaia’s limited precision. In any case, cluster dispersal alone is unlikely to explain expansion on association scales, because it operates on a crossing timescale (Goodwin & Bastian, 2006). For a characteristic vertical size of $\sim$30 pc and a velocity dispersion of $\sim$2.4 km s-1 in this direction, the crossing time is $\gtrsim$10 Myr, greater than the ages of most members.
Alternatively, a large-scale expansion pattern can be imprinted during star formation in a sequential scenario (Elmegreen & Lada, 1977). The age difference between the OB and YSO populations, together with Ser OB2’s geometry, suggests star formation in a cloud accelerated away from the Galactic plane by the expansion of Sh 2-54. We have shown that the overpressurisation of the H ii region is sufficient to drive the required acceleration and to produce a velocity gradient comparable to that observed (Section 7). The expansion and divergence patterns seen in young stellar populations therefore need not arise solely from gas expulsion, but may reflect the initial kinematics of stars upon formation.
@@ Acknowledgements
This research used the NASA/IPAC Infrared Science Archive (IRSA), funded by NASA and operated by Caltech; archival data from the Spitzer Space Telescope, operated by JPL under a NASA contract; and data from ESA’s Gaia mission, processed by the Gaia Data Processing and Analysis Consortium (DPAC), with funding from national institutions, particularly those participating in the Gaia Multilateral Agreement. This work made extensive use of SAOImage DS9 (Joye & Mandel, 2003), the CDS services (Wenger et al., 2000; Ochsenbein et al., 2000; Bonnarel et al., 2000), topcat/stilts (Taylor, 2005, 2025), and the R language (R Core Team, 2022). This work used NASA’s Astrophysics Data System (ADS) under Cooperative Agreement 80NSSC25M7105 and NASA’s Science Explorer (SciX) under Cooperative Agreement 80NSSC21M00561(Kurtz et al., 2000; Grant, 2023).

### v2

@@ Abstract
The Serpens OB2 association ($\ell\sim 18\aas@@fstack{\circ}5$, $b\sim 1\aas@@fstack{\circ}9$, $d=1950\pm 30$ pc) is a large star-forming complex $\sim$65 pc above the Galactic midplane, with a clumpy, elongated structure extending $\sim$50 pc parallel to the plane. We analyse probable association members, including OB stars and low-to-intermediate-mass young stellar objects (YSOs) from the SPICY catalogue. We use 13CO MWISP data to trace the molecular clouds. The OB stars are concentrated toward the centre of the association, coincident with a gap in the molecular clouds, and toward the side nearest the Galactic plane. The YSOs are distributed throughout the association, but cluster around molecular-cloud clumps. Using Gaia DR3 proper motions to probe the association’s internal kinematics, we find aligned stellar velocities on length scales $\lesssim$2 pc, two-point statistics that show increasing velocity differences and predominantly divergent motions at larger separations, and distinct velocities for star clusters within the association. Finally, the association exhibits gradual but statistically significant global expansion perpendicular to the Galactic plane, with a spatial gradient of $0.10\pm 0.02$ km s-1 pc-1. The clumpy stellar distribution, correlated velocities on small scales, and increasingly divergent motions on larger scales are consistent with an initial velocity field inherited from a turbulent molecular cloud modified by stellar feedback. The global vertical expansion may arise from large-scale turbulence or feedback-driven shell expansion, with the H ii region Sh 2-54 preferentially pushing the molecular gas away from the Galactic plane. Ser OB2 demonstrates that the multi-scale expansion of an OB association can begin even while star formation is still ongoing throughout the complex.
@@ 3.2 Member Properties
The $J$ versus JHJ-H colour–magnitude diagram is shown in Fig. 3. Using PARSEC stellar evolution models (Bressan et al., 2012) with an age of 4.4 Myr (Kharchenko et al., 2005), an extinction of $A_{V}\sim 4$ mag reproduces the OB locus but lies significantly bluer than most YSOs. For the 4.4 Myr isochrone to pass through the middle of the YSO distribution on the diagram, it would need to be reddened by $A_{V}\sim 10$ mag. Such an extinction is implausibly high given that many of these stars have Gaia astrometry. For example, a 2 M star with this age and extinction would have a Gaia magnitude of $G=20.6$ mag, which is too faint for reliable astrometry, yet dereddening on the near-infrared diagram suggests that some YSOs are in this mass range. A younger isochrone alleviates this tension. With a 1 Myr isochrone, an extinction of $A_{V}\sim 6$ mag is sufficient to explain the YSO colours, and a 2 M star would then have $G=18.4$ mag, within the range where Gaia provides astrometry. Owing to differential extinction in massive star-forming regions and degeneracies between effective temperature and extinction in spectral-energy-distribution fitting, it is difficult to obtain precise masses and ages for individual stars. Nevertheless, the constraints provided by the colour-magnitude diagram suggest that the YSOs are consistent with an age of $\sim$1 Myr and they are likely to be, on average, younger and more extincted than the OB population.
@@ 3.3 Extrapolated Total Stellar Membership
Assuming our OB member list is complete and finding no evidence of past supernovae, we can extrapolate the total stellar population of the association by scaling from the initial mass function (IMF). The Maschberger (2013) IMF predicts that OB stars ($M>8$ $M_{\odot}$) comprise $\sim$0.7% of all newly formed stellar-mass objects ($M>0.08$ $M_{\odot}$). Given $\sim$80 probable OB stars in Ser OB2 after removing those with inconsistent parallaxes, the association is likely to have a total population of $\sim\!\!10^{4}$ stellar-mass members. For comparison, Luhman (2022) estimates 8300 stellar-mass members in Sco–Cen, meaning that these two associations have similar stellar population numbers.
@@ 6 Molecular Cloud
In the centre of the association, a molecular-gas-free channel extends vertically (in Galactic coordinates) through the cloud. The channel is also visible as a vertical gap in the nebulosity in the Spitzer/IRAC images (Fig. 1), with walls traced by polycyclic aromatic hydrocarbon (PAH) emission. Several OB stars lie in projection within this channel, suggesting they helped ionise the H ii region and clear a passage through the cloud. Because many of the OB stars (including those in NGC 6604) lie below the cloud in Galactic latitude while the thermal chimney extends upward, this channel may have served as the pathway where the H ii region broke through the cloud and inflated the chimney.
Figure 10: The MWISP 13CO J=1J=1–0 zeroth moment map between 25 and 35 km s-1. The greyscale image shows integrated emission in units of $\mathrm{K}\cdot\mathrm{km}$ s-1. OB stars are marked by blue crosses, Ser OB2 clusters (Table 2) are labelled, and the stellar surface density contours from Fig. 4 are drawn in yellow. The molecular-gas-free channel through the centre of the association, also visible in the Spitzer image, is labelled.
@@ 7 Discussion
We observe distinct velocity patterns on different spatial scales, and turbulence provides a natural framework for interpreting this structure. In a turbulent molecular cloud, velocity differences increase with spatial scale (Mac Low & Klessen, 2004), which could be reflected in the velocity statistics of the stars they form (Arnold et al., 2020). On the smallest scales probed by our sample ($\lesssim 2$ pc), the $\theta_{ij}$ statistic indicates that nearby stars have preferentially aligned velocity directions. These scales encompass typical sizes of individual young clusters and subclusters (Lada & Lada, 2003; Kuhn et al., 2014, 2017), whose stars may have formed in the same molecular clump with similar initial velocities. The VSAT statistics probe pairwise velocity differences over more than an order of magnitude in scale. For separations of $\sim$2–20 pc, both statistics show increasing velocity differences with separation, reminiscent of the Larson (1981) relations where cloud velocity dispersions increase with length scale. The VSAT $\Delta v_{ij,D}$ statistic has a positive median for separations >>2 pc, indicating net divergent motions, with divergence increasing with spatial separation. On intermediate scales, several clusters are moving outward, but others deviate from this trend, implying a more complex velocity field. On the largest scales, we detect gradual global expansion of the association along the axis perpendicular to the Galactic plane.
It is plausible that the asymmetric velocity gradient could result from stochastic turbulent fluctuations in a globally unbound system. Nevertheless, it is notable that the gradient is oriented perpendicular to the Galactic plane, where the relative placement of the OB stars, molecular gas, and the thermal chimney provides a natural direction. This feedback geometry may induce an asymmetric expansion of the cloud which could be imprinted on the stars via sequential star formation.
Previous work has shown that the $\sim$4–5 Myr-old OB population in Ser OB2 is energetically capable of driving the large H ii region Sh 2-54 and the associated ‘Stockert chimney’ (Kundt & Mueller, 1987; Forbes, 2000). [^6] An expanding bubble model has been used to successfully describe another thermal chimney in the W4 region, obtaining an expansion age of 2.5 Myr (Basu et al., 1999). While both chimneys are of similar length, the Ser OB2 chimney is much narrower and its ionising source is less certain (Normandeau & Basu, 1999). Although the interface between the Sh 2-54 H ii region and the molecular cloud is likely to be complex (Section 6), the OB stars are preferentially nearer the Galactic plane, so the net force of the expanding H ii region is likely to be directed upward (i.e., in the $\eta$ direction). If the expansion of the Sh 2-54 H ii region has been pushing the cloud away from the midplane, stars forming in this accelerated gas would inherit upward velocities, producing a vertical velocity gradient.
@@ 8 Conclusion
Ser OB2 is a large OB association notable for its youth ($\lesssim 5$ Myr), its moderate height above the Galactic midplane (65 pc), and its role in launching a $\sim$200 pc thermal chimney. We have analysed a sample of Ser OB2 members assembled from OB stars in the ALS III catalogue and YSOs from the SPICY catalogue, with membership vetted using Gaia DR3 parallaxes. Owing to the inclusion of YSOs, this is the first study of the region to incorporate a substantial sample of low- and intermediate-mass members. Although spatial distributions differ moderately, the OB stars and YSOs show similar kinematics, with increasingly divergent motions at larger separations.
@@
  • While previous work indicated an age of 4–5 Myr for the OB population (e.g. Forbes, 2000), the same age would imply implausibly large extinctions for the low-mass YSOs. The YSOs are more consistent with an age of $\sim$1 Myr, implying an overall age spread of a few Myr in the association (Section 3.2).
  • The Ser OB2 members exhibit a clumpy spatial distribution elongated parallel to the Galactic midplane. Although the spatial distributions of OB stars and YSOs overlap, the OB stars are preferentially closer to the Galactic midplane and more concentrated toward the centre of the association in Galactic longitude (Section 4). In this central region, which contains many of the O stars, we identify a gap in the molecular clouds that may serve as a channel for launching the thermal chimney (Section 6).
  • We identified six discrete stellar clusters (NGC 6604, Gum 85, Dobashi 782, 815, 822, and G18.7+1.8) superimposed on a relatively smooth, centrally concentrated stellar distribution (Section 4). Five of these clusters are associated with distinct molecular cloud clumps in 13CO J=1J=1–0 MWISP data cubes.
@@
The disruption of clusters within Ser OB2 may contribute to the region’s overall expansion. However, cluster dispersal alone is unlikely to explain expansion on association scales, because it operates on a crossing timescale (Goodwin & Bastian, 2006). For a characteristic vertical size of $\sim$30 pc and a velocity dispersion of $\sim$2.4 km s-1 in this direction, the crossing time is $\gtrsim$10 Myr, greater than the ages of most members.
Alternatively, the large-scale expansion pattern may be an imprint of the velocity field in the molecular cloud from which the stars formed. Large-scale velocity patterns in the gas may arise from turbulent fluctuations (Mac Low & Klessen, 2004) or from the expansion of a molecular-cloud shell driven by feedback (Elmegreen & Lada, 1977). In Ser OB2, the configuration of the H ii region and molecular cloud would favour expansion away from the Galactic plane, in the same orientation as the association-scale stellar velocity gradient. Furthermore, the difference in the mean ages of the OB stars and YSOs is consistent with the sequential star-formation scenario of Elmegreen & Lada (1977). These results suggest that the expansion and divergence patterns seen in young stellar populations need not arise solely from gas expulsion, but may reflect the initial kinematics of stars upon formation.
@@ Acknowledgements
We thank the referee for insightful perspectives that helped strengthen this work. This research used the NASA/IPAC Infrared Science Archive (IRSA), funded by NASA and operated by Caltech; archival data from the Spitzer Space Telescope, operated by JPL under a NASA contract; and data from ESA’s Gaia mission, processed by the Gaia Data Processing and Analysis Consortium (DPAC), with funding from national institutions, particularly those participating in the Gaia Multilateral Agreement. This work made extensive use of SAOImage DS9 (Joye & Mandel, 2003), the CDS services (Wenger et al., 2000; Ochsenbein et al., 2000; Bonnarel et al., 2000), topcat/stilts (Taylor, 2005, 2025), and the R language (R Core Team, 2022). This work used NASA’s Astrophysics Data System (ADS) under Cooperative Agreement 80NSSC25M7105 and NASA’s Science Explorer (SciX) under Cooperative Agreement 80NSSC21M00561(Kurtz et al., 2000; Grant, 2023).